### 18 Dec 2023 v2

# Preserving feasible space variational quantum algorithm for solving the uncapacitated facility location problem

Abstract
The Uncapacitated Facility Location Problem (UFLP) is an important NP-hard problem with wide applications in many fields, which can be transformed into a constrained optimization problem with unconstrained variables, and we call such problem as Unconstrained-Variables Problem (UVP). The Quantum Alternating Operator Ansatz (QAOA+) is a kind of hybrid quantum-classical algorithm, which can be used to solve the UVP. However, we find that the success probability of QAOA+ may be decreased by the lack of entanglement gates as applied to UVP. In this paper, taking the UFLP as an example, the Preserving Feasible Space-Variational Quantum Algorithm (PFS-VQA) suitable for the UVP was designed. As the mixed operators in QAOA+ preserve the feasible subspace and Hardware-Efficient Ansatz (HEA) reduces the circuit depth, PFS-VQA combines the advantages of both by performing mixed operators on constrained variables and HEA on unconstrained variables. By introducing more CNOT gates and parameters of HEA, PFS-VQA can traverse enough quantum states, thereby improving the success probability. Moreover, since the mixed operators and HEA of PFS-VQA act on different qubits respectively, parallelization can be realized, leading to a lower circuit depth. Finally, the numerical results demonstrate that PFS-VQA decreases the circuit depth, enhances the success probability, and converges faster compared to QAOA+, Quantum Approximation Optimization Algorithm (QAOA), and HEA. The proposed algorithm is flexible as HEA can be replaced if a more efficient ansatz is available. Moreover, our algorithm is general and instructive for solving such UVPs.
pacs:
Valid PACS appear here
preprint: APS/123-QED
I Introduction
By harnessing quantum effects, quantum computers offer computational advantages over classical computers, delivering polynomial or even exponential speedups for specific problems, such as integer factorization 1 , unstructured data search 2 , linear regression 4 ; 5 , quantum error correction 3 , matrix computation 11 ; 12 ; 13 ; 14 , and cryptanalysis 16 . However, the current quantum hardware devices only support a limited number of physical qubits and limited gate fidelity, which makes the above quantum algorithms unable to be implemented on near-term devices.
Variational Quantum Algorithms (VQAs) are a hot class of hybrid quantum-classical algorithms, with one of the core of ansatz design, which are expected to realize quantum advantages on Noisy Intermediate-Scale Quantum (NISQ) devices 17 ; 18 . Ansatz is generally divided into two types: problem-agnostic ansatz and problem-inspired ansatz. The structure of the problem-agnostic ansatz carries no information about the problem itself and is mostly suited for optimization problems, such as the Hardware-Efficient Ansatz (HEA) Hardware-efficient , which has the advantage of reducing the circuit depth as much as possible and being able to be implemented efficiently on a quantum chip. The specific structure of the problem-inspired ansatz typically depends on the task at hand, such as Quantum Approximation Optimization Algorithm (QAOA) QAOA and Quantum Alternating Operator Ansatz (QAOA+) 31 to solve the combinatorial optimization problems, which have also been applied to many problems maxcut2018 ; maxcut2021 ; mvcp2022 ; cc2022 ; 33 ; 34 ; 35 ; 36 ; 37 . HEA and QAOA need to search for the target solution in the Overall Hilbert Space, which may result in an invalid solution. Different from them, the mixed operators of QAOA+ preserves the feasible subspace, allowing the algorithm to search for target solutions in the Feasible Space (a subspace of the entire Hilbert space) of the constrained optimization problems 31 , resulting in the complete elimination of an invalid solution. Since QAOA+ has a significantly smaller search space, finding a solution is potentially more efficient.
The UFLP Kuehn and Hamburger is a combinatorial optimization problem, which is one of the most important NP-hard problems in location theory Daskin2003 , with important applications in fields such as resource allocation Armas and Juan2018 , logistics and transportation Cura2010 , computer networks Lazic2010 , and computer vision Lazic2009 . However, there still exists a research gap in the field of quantum algorithms for addressing this particular issue. Intuitively, the QAOA+ seem to be a good choice for solving UFLP. Nevertheless, we find that the phase-separation operators of QAOA+ introduce lot of CNOT and $R_{Z}$ gates, thereby leading to increased circuit depth. Wang et al 34 have already pointed out that the phase-separation operators only change the phase, and numerical experiments demonstrated that discarding it hardly decreases the success probability. If the UFLP is solved using QAOA+, to facilitate the construction the mixed Hamiltonian, it needs to be transformed into a constrained optimization problem containing unconstrained variables (variables not present in the constraints), and we call it the Unconstrained-Variables Problem (UVP). For UVP, after discarding the phase-separation operator, only single-qubit rotating gates act on the unconstrained variables, lacking entanglement gates, resulting in the inability to traverse enough quantum states, which may lower the success probability of finding a solution. Therefore, QAOA+ is not suitable for such problems. How to design quantum algorithm with low circuit depth to solve UFLP efficiently is particularly important.
As the mixed operators in QAOA+ preserve the feasible subspace and HEA reduces the circuit depth, we consider combining the advantages of both to design a quantum algorithm for UFLP. In this paper, we propose the Preserving Feasible Space-Variational Quantum Algorithm (PFS-VQA), which performs mixed operators and HEA act on different qubits respectively. Specifically, to address UFLP, it is transformed into a UVP in two steps. First, we transform the inequality constraint into equality constraint by introducing slack variables. Second, to facilitate the construction of a mixed Hamiltonian, a penalty function approach is adopted by including a constraint as part of the objective function, resulting in a UVP, where the feasible space is composed of bit strings with a fixed Hamming weight 1. To verify the efficiency of the algorithm, we simulate it for UFLP with 12 instances using MindQuantum 51 , showing that it has at least 54% higher success probability compared to QAOA+, QAOA, and HEA, and at least 75% lower circuit depth compared to QAOA+ and QAOA. Moreover, our algorithm reduces the number of CNOT gates and parameter gates by at least 59% and 53% compared to QAOA+ and QAOA, respectively.
This paper is organized as follows. In Sec. II, the PFS-VQA is proposed. In Sec. III, we apply PFS-VQA to solve UFLP. In Sec. IV, numerical results and analysis are given. Finally, the conclusion is given in Sec. V.
II PFS-VQA
In QAOA+ 31 , the variational ansatz consists of $p$ layers, each containing a phase separator Hamiltonian $H_{P}$ and a mixer Hamiltonian $H_{M}$:
$|\psi_{p}(\overrightarrow{\gamma},\overrightarrow{\beta})\rangle=U(H_{M},\beta_{p})U(H_{P},\gamma_{p})\cdots U(H_{P},\gamma_{1})|x\rangle,$ (1)
where $U(H_{M},\beta_{j})=e^{-i\beta_{j}H_{M}}$, $U(H_{P},\gamma_{j})=e^{-i\gamma_{j}H_{P}}$, $j=1,2,\cdots,p$, and the initial state $|x\rangle$ is a trivial feasible solution, or the uniform superposition state of trivial feasible solutions, and $\overrightarrow{\gamma}=(\gamma_{1},\gamma_{2},\cdots,\gamma_{p})$ and $\overrightarrow{\beta}=(\beta_{1},\beta_{2},\cdots,\beta_{p})$ are variational parameters sets. The variational parameters are optimized on classical computer with the goal of finding the optimal parameters $(\overrightarrow{\gamma^{*}},\overrightarrow{\beta^{*}})$, which are obtained by minimizing the expected value of the phase separator Hamiltonian $H_{P}$
$(\overrightarrow{\gamma^{*}},\overrightarrow{\beta^{*}})\ =arg\ \mathop{min}\limits_{\overrightarrow{\gamma},\overrightarrow{\beta}}\ F_{p}(\overrightarrow{\gamma},\overrightarrow{\beta}),$ (2)
where $F_{p}(\overrightarrow{\gamma},\overrightarrow{\beta})=\langle\psi_{p}(\overrightarrow{\gamma},\overrightarrow{\beta})|H_{P}|\psi_{p}(\overrightarrow{\gamma},\overrightarrow{\beta})\rangle.$
A family of mixing operators $U(H_{M},\beta)$ of QAOA+ limits the state of the system to the feasible space, and results in zero probability of obtaining invalid solutions, which is the advantage of the algorithm. However, a family of phase-separation operators $U(H_{P},\gamma)$ of QAOA+ only changes the phase and has no significant effect on the result 34 . Moreover, since $U(H_{P},\gamma)$ always introduces a lot of CNOT and $R_{Z}$ gates, leading to increased circuit depth, consider discarding $U(H_{P},\gamma)$. HEA has the advantage of reducing the circuit depth as much as possible and being easy to implement efficiently on a quantum chip Hardware-efficient . For UVPs, to reduce the circuit depth, we consider performing HEA Hardware-efficient on non-constrained variables.
In this work, we replace the $U(H_{P},\gamma)$ with the unitary operator $U_{HEA}(\gamma)$ of HEA
$|\psi_{p}(\overrightarrow{\gamma},\overrightarrow{\beta})\rangle=U(H_{M},\beta_{p})U_{HEA}(\overrightarrow{\gamma_{p}})\cdots U_{HEA}(\overrightarrow{\gamma_{1}})|x\rangle,$ (3)
where $\overrightarrow{\gamma_{i}}=(\gamma_{i}^{1},\gamma_{i}^{2},\cdots,\gamma_{i}^{l})$, $i=1,2,\cdots,p$, and $l$ is the number of non-constrained variables. $U(H_{M},\beta_{i})$ acts on variables of the constraints, and $U_{HEA}(\overrightarrow{\gamma_{i}})$ acts on non-constrained variables in constrained optimization problems. The crucial points of PFS-VQA are the initial state $|x\rangle$, the mixing operators $U(H_{M},\beta)$, the unitary operators $U_{HEA}(\gamma)$, and the phase separator Hamiltonian $H_{P}$. The framework of the PFS-VQA is shown in Figure  1.
Figure 1: The framework of the PFS-VQA. The variational parameters were optimized on a classical computer, and a quantum computer was used to evaluate the expected value of the objective function. The mixing operator $U(H_{M},\beta_{i})$ acts on variables of the constraints, and the unitary operator $U_{HEA}(\overrightarrow{\gamma_{i}})$ of HEA acts on non-constrained variables in the constrained optimization problems.
To evaluate the quality of the solution for the optimization problem, we define the success probability as the probability of finding the optimal solution
$P_{success}=|\langle x_{sol}|\psi_{p}(\overrightarrow{\gamma},\overrightarrow{\beta})\rangle|^{2},$ (4)
where $x_{sol}$ is an optimal solution to the problem.
III Apply PFS-VQA to solve UFLP
The UFLP is a combinatorial optimization problem that can be transformed into a variant of UVP. In this section, we first introduce the definition and the mathematical model of UFLP. Then, we solve UFLP using PFS-VQA.
III.1 UFLP
The definition of UFLP is generally described as Zhang2023 : giving the set $K=\{k_{1},\cdots,k_{m}\}$ of customers, where $m$ is the number of customers and $k_{i}$ is the $i$th customer. Giving the set $S=\{s_{1},\cdots,s_{n}\}$ of potential facilities that can be opened, where $n$ is the number of facilities and $s_{j}$ is the $j$th facility. Giving an $m\times n$ matrix $D=\left[d_{ij}\right]_{m\times n}$, where $d_{ij}$ represents the service cost when the $i$th customer receives the service from the $j$th facility. $G=\{g_{1},\cdots,g_{n}\}$ is the set of fixed open cost of facilities, where $g_{j}$ represents the opening cost required by the opening of the $j$th facility. It is worthy to note that the demand of any customer is fulfill by only one facility. The goal of UFLP is to find a set of open facilities and a reasonable allocation scheme between facilities and customers, so that the sum of total service costs and total open facility costs is minimized.
The mathematical model of UFLP Zhang2023 is described as follows
$min\quad\sum_{i=1}^{m}\sum_{j=1}^{n}d_{ij}y_{ij}+\sum_{j=1}^{n}g_{j}x_{j},$ (5)
$s.t.\quad\sum_{j=1}^{n}y_{ij}=1,\quad i=1,2,\cdots,m,$ (6)
$\quad\quad y_{ij}\leq x_{j},\quad i=1,2,\cdots,m,j=1,2,\cdots,n,$ (7)
$\quad\quad y_{ij},x_{j}\in\{0,1\},$ (8)
where $y_{ij}=1$ if $i$th customer gets service from $j$th facility; otherwise $y_{ij}=0$, and $x_{j}=1$ if $j$th facility is opened; otherwise $x_{j}=0$. The number of binary decision variables is $mn+n$. The first term in the Eq. (5) denotes the total service cost, and the second term denotes the total opening cost of the opened facilities. The Eq. (6) ensures that every customer is served by exactly one facility. The Eq. (7) ensures that a customer can be served from a facility only if a facility is opened.
To facilitate solving UFLP, we transformed the above optimization problem into the following standard form by introducing slack variables $z_{ij}$
$min\quad\sum_{i=1}^{m}\sum_{j=1}^{n}d_{ij}y_{ij}+\sum_{j=1}^{n}g_{j}x_{j},$ (9)
$s.t.\quad\sum_{j=1}^{n}y_{ij}=1,\quad i=1,2,\cdots,m,$ (10)
$\quad\quad y_{ij}+z_{ij}-x_{j}=0,\quad i=1,\cdots,m,j=1,\cdots,n,$ (11)
$\quad\quad z_{ij},y_{ij},x_{j}\in\{0,1\}.$ (12)
The number of binary decision variables is $2mn+n$.
PFS-VQA combines the advantages of QAOA+ and HEA by performing mixed operators on constrained variables and HEA on unconstrained variables respectively, which searches for a target solution in the feasible space and also decreases the circuit depth. However, one of the cores of this algorithm is the construction of mixed Hamiltonian.
To facilitate the creation of the mixed Hamiltonian, we adopt the penalty function approach by making the constraint Eq. (11) a part of the objective function Eq. (9), resulting in the following new optimization problem
$min\quad\sum_{i=1}^{m}\sum_{j=1}^{n}d_{ij}y_{ij}+\sum_{j=1}^{n}g_{j}x_{j}+\lambda h(x,y,z),$ (13)
$s.t.\quad\sum_{j=1}^{n}y_{ij}=1,\quad i=1,2,\cdots,m,$ (14)
$\quad\quad z_{ij},y_{ij},x_{j}\in\{0,1\},$ (15)
where $h(x,y,z)=\sum_{i=1}^{m}\sum_{j=1}^{n}(y_{ij}+z_{ij}-x_{j})^{2}$, $\lambda$ is the penalty, determined empirically. The number of constraint variables in Eq. (14) is $mn$, and the number of non-constrained variables in Eq. (13) is $mn+n$. Next, we will solve the UFLP via PFS-VQA designed in section  II.
III.2 PFS-VQA FOR UFLP
The crucial points of PFS-VQA are the initial state $|x\rangle$, the mixing operators $U(H_{M},\beta)$, the unitary operators $U_{HEA}(\gamma)$, and the phase separator Hamiltonian $H_{P}$ for UFLP. For initial state $|x\rangle$, according to Eq. (14), the state $\overbrace{|10\cdots 0_{n-1}}^{n}\overbrace{10\cdots 0_{2n-1}}^{n}\cdots\overbrace{10\cdots 0_{mn-1}}^{n}\overbrace{0\cdots 0_{2mn+n-1}}^{mn+n}\rangle$ can be obtained, which is a trivial feasible solution.
The mixing operators $U(H_{M},\beta)=e^{-i\beta H_{M}}$ depend on Eq. (14) and its structure, and its core is to construct $H_{M}$. To maintain the Hamming weight 1 of the constraint Eq. (14), the mixing Hamiltonian $H_{M}$ is expressed as follows 31 ; 33
$H_{M}=\sum_{i=0}^{m-1}\sum_{j=0}^{n-2}X_{j+i\cdot n}X_{j+i\cdot n+1}+Y_{j+i\cdot n}Y_{j+i\cdot n+1},$ (16)
where $X$, and $Y$ represents Pauli-$X$ operation, and Pauli-$Y$ operation respectively.
The unitary operators $U_{HEA}(\gamma)$ of HEA consist of single-qubit rotating gates and entangled gates as shown in Figure  1. The phase separator Hamiltonian $H_{P}$ is obtained by replacing binary variables $x$, $y$, $z$ in Eq. (13) with $\frac{I-Z}{2}$
$H_{P}$ $=\sum_{i=1}^{m}\sum_{j=1}^{n}d_{ij}\frac{I-Z_{n*(i-1)+j-1}}{2}+\sum_{j=1}^{n}g_{j}\frac{I-Z_{m*n+j-1}}{2}$
$+\lambda\sum_{i=1}^{m}\sum_{j=1}^{n}(\frac{I-Z_{n*(i-1)+j-1}}{2}+\frac{I-Z_{n*(m+i)+j-1}}{2}$
$-\frac{I-Z_{m*n+j-1}}{2})^{2},$ (17)
where $Z$ represents Pauli-$Z$ operation and the subscript represents the qubit of action. After the three key points are structured, we apply PFS-VQA to solve the UFLP and perform numerical simulation experiments.
IV Numerical Simulation
In this section, we perform numerical experiments using the MindSpore Quantum 51 . We study 12 instances for three different problem sizes of the UFLP given in Table  1 to benchmark the performance of PFS-VQA. Details of the instances are given in Appendix A, corresponding to 10, 14, and 22 qubits, respectively. To find the optimal variational parameters, an Adam optimizer Adam2015 is utilized which is updated with gradients by an adaptive moment estimation algorithm. For the selection of initial parameters, the random initialization method is adopted.
Table 1: The instances of UFLP.
$m\times n$
Qubit
Instances
Optimal Value
2 $\times$ 2 10
Instance 1
16
Instance 2
42
Instance 3
30
Instance 4
39
Instance 5
52
3 $\times$ 2 14
Instance 6
21
Instance 7
42
Instance 8
40
Instance 9
35
Instance 10
43
5 $\times$ 2 22
Instance 11
82
Instance 12
95
Figure 2: Comparison of the performance of standard QAOA, QAOA+, HEA, and PFS-VQA for the UFLP with 12 instances. The circuit depth and the average success probability as a function of the layers are shown for the standard QAOA, QAOA+, HEA, and PFS-VQA. (a)-(f) The optimization trajectories from optimizer Adam are plotted versus the number of circuit evaluations. For both cases $m=2$, $n=2$ and $m=3$, $n=2$, we perform numerical simulations up to $p=6$ using random initialization. However, owing to the large number of qubits and high circuit depth involved in case $m=5$, $n=2$, only simulations from $p=1$ to $p=3$ are performed. (a)-(b) $m=2$, $n=2$, (c)-(d) $m=3$, $n=2$, (e)-(f) $m=5$, $n=2$.
To verify the high efficiency of PFS-VQA, we also apply QAOA, QAOA+, and HEA to UFLP and compared them, giving the Hamiltonians and the corresponding circuits, as specified in Appendix B, C, and D respectively. In Fig. 2, we performed numerical simulations up to $p=6$ using random initialization for 12 instances. The circuit depth and average success probability as a function of the layers are shown for the standard QAOA, QAOA+, HEA, and PFS-VQA, respectively.
Specifically, for $m=2$, $n=2$, in Fig. 2 (a) and (b), PFS-VQA has a higher success probability than the other three algorithms, at least 54% higher. Meanwhile, PFS-VQA has the same circuit depth as HEA, which is at least 75% lower than QAOA and QAOA+. For $m=3$, $n=2$, in Fig. 2 (c) and (d), PFS-VQA has a success probability of at least 58% higher compared to the other three algorithms. Moreover, the circuit depth of PFS-VQA is about 26% lower than HEA, and at least 83% lower than QAOA and QAOA+. For $m=5$, $n=2$, in Fig. 2 (e) and (f), compared to the other three algorithms, PFS-VQA has at least five times higher success probability. And, the circuit depth of PFS-VQA is about 43% lower than HEA, and at least 87% lower than QAOA and QAOA+. We find that the success probability of PFS-VQA is higher than the other three algorithms by an increasing percentage as the number of qubits increases. Directly, we conjecture that the performance of this algorithm is less affected by the increase in qubits than QAOA, QAOA+ and HEA.
The results show that PFS-VQA achieves a significantly higher success probability at low circuit depths compared to the QAOA, QAOA+, and HEA. One of the keys to these four algorithms is to construct the phase separator Hamiltonian. In QAOA and HEA, for constrained optimization problems, the common method is to incorporate hard constraints into the target function as a penalty item, and then convert the target function into a phase separator Hamiltonian 30 . These two algorithms need to search for the target solutions in the Overall Hilbert Space 31 . Differently, PFS-VQA encodes the constraints directly into quantum circuits, limiting the state of the system to the Feasible Space (a subspace of the entire Hilbert space) of the constrained optimization problems, which implies a higher probability of searching a solution compared to these two algorithms. Furthermore, for the Eq. (13) with non-constrained variables, QAOA+ performs single-qubit rotating gates $R_{X}$ and the phase-separation operators $U(H_{P},\gamma)$ for non-constrained variables (see Appendix C for details). Nevertheless, the $U(H_{P},\gamma)$ of QAOA+ introduce lot of CNOT and $R_{Z}$ gates, thereby leading to increased circuit depth. Wang et al 34 have already pointed out that the $U(H_{P},\gamma)$ only change the phase, and numerical experiments demonstrated that discarding it hardly decreases the success probability. For unconstrained variables, after discarding the phase-separation operator, only single-qubit rotating gates act on the unconstrained variables, lacking entanglement gates, resulting in the inability to traverse enough quantum states, which may lower the success probability of finding a solution. Different from QAOA+, PFS-VQA performs the mixed operators and HEA act on different qubits, parallelization can be realized, leading to a lower circuit depth. Therefore, PFS-VQA has lower circuit depth and higher success probability at the same low-layer $p$ compared to the standard QAOA, QAOA+, and HEA.
Figure 3: The comparison of the iterations.
To evaluate the convergence speed of our algorithm, we apply QAOA+, QAOA, and HEA to UFLP and compared them. We selected the largest scale instance and plotted Fig. 3 through numerical simulation. The loss value as a function of the iterations with $p=3$ is shown for the standard QAOA, QAOA+, HEA, and PFS-VQA, respectively. The results show that our algorithm requires only about 125 iterations to converge to the optimal value, which is faster and better convergence compared to the other three algorithms.
Figure 4: Resource comparison of the standard QAOA, QAOA+, HEA, and PFS-VQA for the UFLP with three different sizes. (a), (b), and (c) show the comparison of four algorithms for the 10, 14, and 22 qubits, respectively.
A crucial question, especially for near-term devices, is how the different algorithms compare with respect to resource overhead. We compare quantum resources in terms of the number of CNOT gates, parameter gates, and parameters. Fig. 4 shows the number of parameters, CNOTs, and parameter gates for the four algorithms with $p=2$. The results show that our algorithm reduces both the number of CNOT gates by at least 53% and parameter gates by at least 59% compared to QAOA and QAOA+, and the number of parameters by at least 33% and parameter gates by at least 13% compared to HEA. However, our algorithm has significantly more parameters than QAOA and QAOA+, and more CNOT gates than HEA. In other words, the PFS-VQA obtains better performance in UFLP at the cost of introducing more variational parameters and CNOTs.
V CONCLUSION
In conclusion, we proposed PFS-VQA for solving such UVPs, a combined ansatz that incorporates the advantages of HEA and QAOA+. To verify the high efficiency of PFS-VQA, we also applied QAOA, QAOA+, and HEA to UFLP, as well as compare the performance of these four algorithms. The Hamiltonians and the corresponding circuits are given as specified in Appendix B, C, and D respectively. We tested the performance of several instances of UFLP with 10, 14, and 22 qubits, details of the instances are given in Appendix A, finding that PFS-VQA always outperforms the other three algorithms (see Fig. 2 for details).
Specifically, for $m=2$, $n=2$, in Fig. 2 (a) and (b), PFS-VQA has a higher success probability than the other three algorithms, at least 54% higher. Meanwhile, PFS-VQA has the same circuit depth as HEA, which is at least 75% lower than QAOA and QAOA+. For $m=3$, $n=2$, in Fig. 2 (c) and (d), PFS-VQA has a success probability of at least 58% higher compared to the other three algorithms. Moreover, the circuit depth of PFS-VQA is about 26% lower than HEA, and at least 83% lower than QAOA and QAOA+. For $m=5$, $n=2$, in Fig. 2 (e) and (f), compared to the other three algorithms, PFS-VQA has at least five times higher success probability. And, the circuit depth of PFS-VQA is about 43% lower than HEA, and at least 87% lower than QAOA and QAOA+. For the largest scale instance, when $p=3$, our algorithm requires only about 125 iterations to converge to the optimal value, which is faster and better convergence compared to the other three algorithms (see Fig. 3 for details). In addition, we also compared the number of CNOT gates, parameter gates, and parameters of these four algorithms, and the results showed that our algorithm reduced the number of CNOT gates and parameter gates compared to QAOA and QAOA+, and both parameters and parameter gates are reduced compared to HEA (see Fig. 4 for details). The proposed algorithm is flexible as HEA can be replaced if a more efficient ansatz is available. Moreover, our algorithm is general and instructive for solving such UVPs.
Acknowledgements.
This work is supported by the Beijing Natural Science Foundation (Grant No. 4222031), the National Natural Science Foundation of China (Grant Nos. 61976024, 61972048, 62272056), and the 111 Project B21049.
Appendix A Details of instances
In section  IV, numerical simulations are performed for 12 instances as shown in Table  2, Table  3, and Table  4.
Table 2: The Instance 1-Instance 5 of UFLP.
$m\times n$
Qubit
The service matrix $D$
The set $G$ of fixed open cost of facilities
Optimal Value
2 $\times$ 2 10
$\begin{bmatrix}6&10\\ 3&5\end{bmatrix}$
$\left\{7,7\right\}$
16
$\begin{bmatrix}16&10\\ 13&15\end{bmatrix}$
$\left\{17,17\right\}$
42
$\begin{bmatrix}8&15\\ 20&15\end{bmatrix}$
$\left\{9,10\right\}$
30
$\begin{bmatrix}6&20\\ 13&25\end{bmatrix}$
$\left\{20,20\right\}$
39
$\begin{bmatrix}25&20\\ 6&17\end{bmatrix}$
$\left\{27,15\right\}$
52
Table 3: The Instance 6-Instance 10 of UFLP.
$m\times n$
Qubit
The service matrix $D$
The set $G$ of fixed open cost of facilities
Optimal Value
3 $\times$ 2 14
$\begin{bmatrix}6&10\\ 3&1\\ 5&4\end{bmatrix}$
$\left\{7,7\right\}$
21
$\begin{bmatrix}16&10\\ 13&5\\ 4&10\end{bmatrix}$
$\left\{17,17\right\}$
42
$\begin{bmatrix}6&10\\ 3&5\\ 4&1\end{bmatrix}$
$\left\{27,27\right\}$
40
$\begin{bmatrix}6&20\\ 3&15\\ 24&1\end{bmatrix}$
$\left\{10,15\right\}$
35
$\begin{bmatrix}56&10\\ 23&5\\ 4&18\end{bmatrix}$
$\left\{27,10\right\}$
43
Table 4: The Instance 11-Instance 12 of UFLP.
$m\times n$
Qubit
The service matrix $D$
The set $G$ of fixed open cost of facilities
Optimal Value
5 $\times$ 2 22
$\begin{bmatrix}16&10\\ 13&15\\ 14&10\\ 15&18\\ 20&25\end{bmatrix}$
$\left\{7,7\right\}$
82
$\begin{bmatrix}16&10\\ 13&15\\ 14&10\\ 15&18\\ 20&25\end{bmatrix}$
$\left\{17,17\right\}$
95
Appendix B QAOA for UFLP
The general framework of QAOA is similar to QAOA+ as shown in section  II, which we will not review here. Similarly, the crucial points of QAOA are the initial state $|x\rangle$, the mixing operators $U(H_{M},\beta)$, and the phase-separation operators $U(H_{P},\gamma)$ for UFLP. The initial state is generally chosen as a uniform superposition state, i.e. $|+\rangle^{\bigotimes N}$, $N=2mn+n$. For the mixing operators $U(H_{M},\beta)=e^{-i\beta H_{M}}$, and its core is to construct $H_{M}$, where $H_{M}$ is set to $\sum_{j=0}^{N-1}X_{j}$. The crucial to the algorithm is the design of the phase separator Hamiltonian, which includes solutions of the UFLP. For the following standard form of the UFLP mathematical model
$min\quad\sum_{i=1}^{m}\sum_{j=1}^{n}d_{ij}y_{ij}+\sum_{j=1}^{n}g_{j}x_{j},$ (18)
$s.t.\quad\sum_{j=1}^{n}y_{ij}=1,\quad i=1,2,\cdots,m,$ (19)
$\quad\quad y_{ij}+z_{ij}-x_{j}=0,\quad i=1,\cdots,m,j=1,\cdots,n,$ (20)
$\quad\quad z_{ij},y_{ij},x_{j}\in\{0,1\}.$ (21)
To construct phase separator Hamiltonian $H_{P}$, we adopt the common method of introducing the hard constraint as a penalty term into the objective function 26 ; 27 ; 28 ; 29 ; 30 . The following unconstrained optimization problem is obtained
$min\quad\sum_{i=1}^{m}\sum_{j=1}^{n}d_{ij}y_{ij}+\sum_{j=1}^{n}g_{j}x_{j}+\lambda(\sum_{i=1}^{m}(\sum_{j=1}^{n}y_{ij}-1)^{2}+\sum_{i=1}^{m}\sum_{j=1}^{n}(y_{ij}+z_{ij}-x_{j})^{2}),$ (22)
where $\lambda$ is the penalty, determined empirically.
The phase separator Hamiltonian $H_{P}$ is obtained by replacing binary variables $x$, $y$, $z$ in Eq. (22) with $\frac{I-Z}{2}$
$H_{P}$ $=\sum_{i=1}^{m}\sum_{j=1}^{n}d_{ij}\frac{I-Z_{n*(i-1)+j-1}}{2}+\sum_{j=1}^{n}g_{j}\frac{I-Z_{m*n+j-1}}{2}$
$+\lambda(\sum_{i=1}^{m}(\sum_{j=1}^{n}\frac{I-Z_{n*(i-1)+j-1}}{2}-I)^{2}+\sum_{i=1}^{m}\sum_{j=1}^{n}(\frac{I-Z_{n*(i-1)+j-1}}{2}+\frac{I-Z_{n*(m+i)+j-1}}{2}-\frac{1-Z_{m*n+j-1}}{2})^{2}),$ (23)
where $Z$ represents Pauli-$Z$ operation and the subscript represents the qubit of action. After the three key points are structured, we apply QAOA to solve the UFLP and perform numerical simulation experiments, and the corresponding 2-layer quantum circuit is shown in Figure  5.
Figure 5: The overall 2-layer circuit of QAOA.
Appendix C QAOA+ for UFLP
The crucial points of QAOA+ are the initial state $|x\rangle$, the mixing operators $U(H_{M},\beta)$, the phase-separation operators $U(H_{P},\gamma)$ for UFLP. For initial state $|x\rangle$, according to Eq. (14), the state $\overbrace{|10\cdots 0_{n-1}}^{n}\overbrace{10\cdots 0_{2n-1}}^{n}\cdots\overbrace{10\cdots 0_{mn-1}}^{n}\overbrace{0\cdots 0_{2mn+n-1}}^{mn+n}\rangle$ can be obtained, which is a trivial feasible solution.
The mixing operators $U(H_{M},\beta)=e^{-i\beta H_{M}}$ depends on Eq. (14) and its structure, and its core is to construct $H_{M}$. To maintain the Hamming weight 1 of the constraint Eq. (14), the mixing Hamiltonian $H_{M}$ is expressed as follows 31 ; 33
$H_{M}=\sum_{i=0}^{m-1}\sum_{j=0}^{n-2}X_{j+i\cdot n}X_{j+i\cdot n+1}+Y_{j+i\cdot n}Y_{j+i\cdot n+1},$ (24)
where $X$, and $Y$ represents Pauli-$X$ operation, and Pauli-$Y$ operation respectively.
The phase-separation operators $U(H_{P},\gamma)$ depends on Eq. (13), and its core is to construct $H_{P}$. The objective function $f=\sum_{i=1}^{m}\sum_{j=1}^{n}d_{ij}y_{ij}+\sum_{j=1}^{n}g_{j}x_{j}+\lambda\sum_{i=1}^{m}\sum_{j=1}^{n}(y_{ij}+z_{ij}-x_{j})^{2}$, and the phase separator Hamiltonian $H_{P}$ is obtained by replacing binary variables $x$, $y$, $z$ in Eq. (13) with $\frac{I-Z}{2}$
$H_{P}$ $=\sum_{i=1}^{m}\sum_{j=1}^{n}d_{ij}\frac{I-Z_{n*(i-1)+j-1}}{2}+\sum_{j=1}^{n}g_{j}\frac{I-Z_{m*n+j-1}}{2}$
$+\lambda\sum_{i=1}^{m}\sum_{j=1}^{n}(\frac{I-Z_{n*(i-1)+j-1}}{2}+\frac{I-Z_{n*(m+i)+j-1}}{2}-\frac{I-Z_{m*n+j-1}}{2})^{2},$ (25)
where $Z$ represents Pauli-$Z$ operation and the subscript represents the qubit of action. After the three key points are structured, we apply QAOA+ to solve the UFLP and perform numerical simulation experiments, and the corresponding 2-layer quantum circuit is shown in Figure  6.
Figure 6: The overall 2-layer circuit of QAOA+.
Appendix D Hardware Efficient Ansatz for UFLP
The Hardware Efficient Ansatz (HEA) Hardware-efficient is a generic name used for ansatzes that are aimed at reducing the circuit depth needed to implement the parametric circuit when using a given quantum hardware. A major advantage of the HEA is its multifunctionality, as it can adapt to encoding symmetries HEA Bryan ; HEA Matthew and bring the relevant qubits closer to depth reduction HEA Nikolay , as well as being particularly useful for investigating Hamiltonians similar to the device’s interactions HEA Christian . Optimizing the variational parameters of the circuit is performed on classical computers, aiming to find the optimal parameters by minimizing the expected value of the phase separator Hamiltonian. For UFLP, similar to QAOA, the phase separator Hamiltonian is given in Eq. (23). For Instance 1-Instance 5, the corresponding 2-layer quantum circuit is shown in Figure  7.
Figure 7: The overall 2-layer quantum circuit of HEA.

### 20 Dec 2023 v3

@@
 However, the current quantum hardware devices only support a limited number of physical qubits and limited gate fidelity, which makes the above quantum algorithms unable to be implemented on near-term devices.
-Variational Quantum Algorithms (VQAs) are a hot class of hybrid quantum-classical algorithms, with one of the core of ansatz design, which are expected to realize quantum advantages on Noisy Intermediate-Scale Quantum (NISQ) devices 17 ; 18 .
+Variational Quantum Algorithms (VQAs) are a hot class of hybrid quantum-classical algorithms, promising to realize quantum advantages on Noisy Intermediate-Scale Quantum (NISQ) devices 17 ; 18 , and one of the core is the design of ansatz.
 Ansatz is generally divided into two types: problem-agnostic ansatz and problem-inspired ansatz.
 The structure of the problem-agnostic ansatz carries no information about the problem itself and is mostly suited for optimization problems, such as the Hardware-Efficient Ansatz (HEA) Hardware-efficient , which has the advantage of reducing the circuit depth as much as possible and being able to be implemented efficiently on a quantum chip.
-The specific structure of the problem-inspired ansatz typically depends on the task at hand, such as Quantum Approximation Optimization Algorithm (QAOA) QAOA and Quantum Alternating Operator Ansatz (QAOA+) 31 to solve the combinatorial optimization problems, which have also been applied to many problems maxcut2018 ; maxcut2021 ; mvcp2022 ; cc2022 ; 33 ; 34 ; 35 ; 36 ; 37 .
+The specific structure of the problem-inspired ansatz typically depends on the task at hand, such as Quantum Approximation Optimization Algorithm (QAOA) QAOA and Quantum Alternating Operator Ansatz (QAOA+) 31 to solve the combinatorial optimization problems, which have been applied to many problems maxcut2018 ; maxcut2021 ; mvcp2022 ; cc2022 ; 33 ; 34 ; 35 ; 36 ; 37 .
 HEA and QAOA need to search for the target solution in the Overall Hilbert Space, which may result in an invalid solution.
@@
 Since QAOA+ has a significantly smaller search space, finding a solution is potentially more efficient.
-The UFLP Kuehn and Hamburger is a combinatorial optimization problem, which is one of the most important NP-hard problems in location theory Daskin2003 , with important applications in fields such as resource allocation Armas and Juan2018 , logistics and transportation Cura2010 , computer networks Lazic2010 , and computer vision Lazic2009 .
+The UFLP Kuehn and Hamburger is a combinatorial optimization problem, which is one of the most important NP-hard problems in location theory Daskin2003 , with important applications in fields such as resource allocation Armas and Juan2018 , logistics and transportation Cura2010 , and computer vision Lazic2009 .
 However, there still exists a research gap in the field of quantum algorithms for addressing this particular issue.
@@
 Second, to facilitate the construction of a mixed Hamiltonian, a penalty function approach is adopted by including a constraint as part of the objective function, resulting in a UVP, where the feasible space is composed of bit strings with a fixed Hamming weight 1.
-To verify the efficiency of the algorithm, we simulate it for UFLP with 12 instances using MindQuantum 51 , showing that it has at least 54% higher success probability compared to QAOA+, QAOA, and HEA, and at least 75% lower circuit depth compared to QAOA+ and QAOA.
-Moreover, our algorithm reduces the number of CNOT gates and parameter gates by at least 59% and 53% compared to QAOA+ and QAOA, respectively.
+To verify the efficiency of the algorithm, we simulate it for UFLP with 12 instances using MindQuantum 51 , showing that it has at least 54% higher success probability and requires only about 125 iterations to converge to the optimum, with faster and better convergence compared to QAOA+, QAOA, and HEA.
+In addition, our algorithm is at least 75% lower circuit depth and reduces the number of CNOT gates and parameter gates by at least 59% and 53% compared to QAOA+ and QAOA, respectively.
 This paper is organized as follows.

### 11 Jan 2024 v4

@@
-Preserving feasible space variational quantum algorithm for solving the uncapacitated facility location problem
+Variational quantum algorithm-preserving feasible space for solving the uncapacitated facility location problem
 Abstract
-The Uncapacitated Facility Location Problem (UFLP) is an important NP-hard problem with wide applications in many fields, which can be transformed into a constrained optimization problem with unconstrained variables, and we call such problem as Unconstrained-Variables Problem (UVP).
-The Quantum Alternating Operator Ansatz (QAOA+) is a kind of hybrid quantum-classical algorithm, which can be used to solve the UVP.
-However, we find that the success probability of QAOA+ may be decreased by the lack of entanglement gates as applied to UVP.
-In this paper, taking the UFLP as an example, the Preserving Feasible Space-Variational Quantum Algorithm (PFS-VQA) suitable for the UVP was designed.
-As the mixed operators in QAOA+ preserve the feasible subspace and Hardware-Efficient Ansatz (HEA) reduces the circuit depth, PFS-VQA combines the advantages of both by performing mixed operators on constrained variables and HEA on unconstrained variables.
-By introducing more CNOT gates and parameters of HEA, PFS-VQA can traverse enough quantum states, thereby improving the success probability.
-Moreover, since the mixed operators and HEA of PFS-VQA act on different qubits respectively, parallelization can be realized, leading to a lower circuit depth.
-Finally, the numerical results demonstrate that PFS-VQA decreases the circuit depth, enhances the success probability, and converges faster compared to QAOA+, Quantum Approximation Optimization Algorithm (QAOA), and HEA.
-The proposed algorithm is flexible as HEA can be replaced if a more efficient ansatz is available.
-Moreover, our algorithm is general and instructive for solving such UVPs.
+The Quantum Alternating Operator Ansatz (QAOA+) is one of the Variational Quantum Algorithm (VQA) specifically developed to tackle combinatorial optimization problems by exploring the feasible space in search of a target solution.
+For constrained optimization problems with unconstrained variables, which we call Unconstrained-Variables Problems (UVPs), the mixed operators in the QAOA+ circuit are applied to the constrained variables, while the single-qubit rotating gates $R_{X}$ operate on the unconstrained variables.
+The expressibility of this circuit is limited by the shortage of two-qubit gates and the parameter sharing in the $R_{X}$, which consequently impacts the performance of QAOA+ for solving UVPs.
+Therefore, it is crucial to develop a suitable ansatz for UVPs.
+In this paper, we propose the Variational Quantum Algorithm-Preserving Feasible Space (VQA-PFS) ansatz, exemplified by the Uncapacitated Facility Location Problem (UFLP), that applies mixed operators on constrained variables while employing Hardware-Efficient Ansatz (HEA) on unconstrained variables.
+The numerical results demonstrate that VQA-PFS significantly enhances the success probability and exhibits faster convergence compared to QAOA+, Quantum Approximation Optimization Algorithm (QAOA), and HEA.
+Furthermore, VQA-PFS reduces the circuit depth dramatically in comparison to QAOA+ and QAOA.
+Our algorithm is general and instructive in tackling UVPs.
 pacs:
@@
 However, the current quantum hardware devices only support a limited number of physical qubits and limited gate fidelity, which makes the above quantum algorithms unable to be implemented on near-term devices.
-Variational Quantum Algorithms (VQAs) are a hot class of hybrid quantum-classical algorithms, promising to realize quantum advantages on Noisy Intermediate-Scale Quantum (NISQ) devices 17 ; 18 , and one of the core is the design of ansatz.
-Ansatz is generally divided into two types: problem-agnostic ansatz and problem-inspired ansatz.
+Variational Quantum Algorithms (VQAs) are a hot class of hybrid quantum-classical algorithms, promising to realize quantum advantages on Noisy Intermediate-Scale Quantum (NISQ) devices 17 ; 18 .
+The ansatz design of VQAs is important, and ansatz is generally divided into two types: problem-agnostic ansatz and problem-inspired ansatz.
 The structure of the problem-agnostic ansatz carries no information about the problem itself and is mostly suited for optimization problems, such as the Hardware-Efficient Ansatz (HEA) Hardware-efficient , which has the advantage of reducing the circuit depth as much as possible and being able to be implemented efficiently on a quantum chip.
@@
 HEA and QAOA need to search for the target solution in the Overall Hilbert Space, which may result in an invalid solution.
-Different from them, the mixed operators of QAOA+ preserves the feasible subspace, allowing the algorithm to search for target solutions in the Feasible Space (a subspace of the entire Hilbert space) of the constrained optimization problems 31 , resulting in the complete elimination of an invalid solution.
-Since QAOA+ has a significantly smaller search space, finding a solution is potentially more efficient.
-The UFLP Kuehn and Hamburger is a combinatorial optimization problem, which is one of the most important NP-hard problems in location theory Daskin2003 , with important applications in fields such as resource allocation Armas and Juan2018 , logistics and transportation Cura2010 , and computer vision Lazic2009 .
-However, there still exists a research gap in the field of quantum algorithms for addressing this particular issue.
-Intuitively, the QAOA+ seem to be a good choice for solving UFLP.
-Nevertheless, we find that the phase-separation operators of QAOA+ introduce lot of CNOT and $R_{Z}$ gates, thereby leading to increased circuit depth.
-Wang et al 34 have already pointed out that the phase-separation operators only change the phase, and numerical experiments demonstrated that discarding it hardly decreases the success probability.
-If the UFLP is solved using QAOA+, to facilitate the construction the mixed Hamiltonian, it needs to be transformed into a constrained optimization problem containing unconstrained variables (variables not present in the constraints), and we call it the Unconstrained-Variables Problem (UVP).
-For UVP, after discarding the phase-separation operator, only single-qubit rotating gates act on the unconstrained variables, lacking entanglement gates, resulting in the inability to traverse enough quantum states, which may lower the success probability of finding a solution.
+QAOA+ can restrict the state of the system to the Feasible Space (a subspace of the entire Hilbert space), resulting in zero probability of obtaining invalid solutions, which implies a prominent advantage compared to QAOA and HEA.
+For constrained optimization problems without unconstrained variables (variables not included in constraints), one-layer QAOA+ circuit is shown in Figure  1 (a), with the mixed operator $e^{-i\beta_{1}H_{M}}$ applied to all qubits.
+For constrained optimization problems with unconstrained variables, one-layer QAOA+ circuit is shown in Figure  1 (b).
+For simplicity, we call such problems as Unconstrained-Variables Problems (UVPs).
+In Figure  1 (b), the mixed operator $e^{-i\beta_{1}H_{M}}$ acts on the constrained variables, while only the single-qubit rotating gates $R_{X}$ act on the unconstrained variables.
+The representation capability of this circuit is limited by the lack of two-qubit gates and the sharing of parameters in the $R_{X}$, which in turn may affect the performance of QAOA+ to solve UVPs.
 Therefore, QAOA+ is not suitable for such problems.
-How to design quantum algorithm with low circuit depth to solve UFLP efficiently is particularly important.
-As the mixed operators in QAOA+ preserve the feasible subspace and HEA reduces the circuit depth, we consider combining the advantages of both to design a quantum algorithm for UFLP.
-In this paper, we propose the Preserving Feasible Space-Variational Quantum Algorithm (PFS-VQA), which performs mixed operators and HEA act on different qubits respectively.
-Specifically, to address UFLP, it is transformed into a UVP in two steps.
+It is particularly important to design a quantum algorithm for solving UVPs efficiently.
+Figure 1: (a) One layer QAOA+ circuit for constrained optimization problems without unconstrained variables.
+(b) One layer QAOA+ circuit for constrained optimization problems including unconstrained variables, where $X_{\beta_{1}}$ represents the single-qubit rotating gate $R_{X}$.
+The Uncapacitated Facility Location Problem (UFLP) Kuehn and Hamburger is a combinatorial optimization problem, one of the most important NP-hard problems.
+However, there still exists a research blank in the field of quantum algorithms for addressing this particular issue.
+The UFLP can be transformed into a UVP in two steps.
 First, we transform the inequality constraint into equality constraint by introducing slack variables.
-Second, to facilitate the construction of a mixed Hamiltonian, a penalty function approach is adopted by including a constraint as part of the objective function, resulting in a UVP, where the feasible space is composed of bit strings with a fixed Hamming weight 1.
+Second, to facilitate the construction of the mixed Hamiltonian, a penalty function approach is adopted by including a constraint as part of the objective function, thus obtaining a UVP where the feasible space is composed of bit strings with a fixed Hamming weight 1.
+In this paper, we propose a novel ansatz, taking the UFLP as an example, that performs mixed operators $e^{-i\beta H_{M}}$ and HEA act on different qubits respectively.
+This ansatz is called the Variational Quantum Algorithm-Preserving Feasible Space (VQA-PFS) since it preserves the feasible space.
 To verify the efficiency of the algorithm, we simulate it for UFLP with 12 instances using MindQuantum 51 , showing that it has at least 54% higher success probability and requires only about 125 iterations to converge to the optimum, with faster and better convergence compared to QAOA+, QAOA, and HEA.
-In addition, our algorithm is at least 75% lower circuit depth and reduces the number of CNOT gates and parameter gates by at least 59% and 53% compared to QAOA+ and QAOA, respectively.
+In addition, the circuit depth of our algorithm is reduced by at least 75%, and the number of CNOT gates and parameter gates are reduced by at least 59% and 53%, respectively, compared to QAOA+ and QAOA.
 This paper is organized as follows.
 In Sec.
-II, the PFS-VQA is proposed.
+II, the VQA-PFS is proposed.
 In Sec.
-III, we apply PFS-VQA to solve UFLP.
+III, we apply VQA-PFS to solve UFLP.
 In Sec.
@@
 V.
-II PFS-VQA
+II VQA-PFS
 In QAOA+ 31 , the variational ansatz consists of $p$ layers, each containing a phase separator Hamiltonian $H_{P}$ and a mixer Hamiltonian $H_{M}$:
@@
 A family of mixing operators $U(H_{M},\beta)$ of QAOA+ limits the state of the system to the feasible space, and results in zero probability of obtaining invalid solutions, which is the advantage of the algorithm.
-However, a family of phase-separation operators $U(H_{P},\gamma)$ of QAOA+ only changes the phase and has no significant effect on the result 34 .
-Moreover, since $U(H_{P},\gamma)$ always introduces a lot of CNOT and $R_{Z}$ gates, leading to increased circuit depth, consider discarding $U(H_{P},\gamma)$.
 HEA has the advantage of reducing the circuit depth as much as possible and being easy to implement efficiently on a quantum chip Hardware-efficient .
-For UVPs, to reduce the circuit depth, we consider performing HEA Hardware-efficient on non-constrained variables.
-In this work, we replace the $U(H_{P},\gamma)$ with the unitary operator $U_{HEA}(\gamma)$ of HEA
+For UVPs, one-layer QAOA+ circuit is shown in Figure  1 (b), with the mixed operator $e^{-i\beta_{1}H_{M}}$ acting on the constrained variables, and only the single-qubit rotating gates $R_{X}$ acting on the unconstrained variables.
+The representation capability of this circuit is limited by the lack of two-qubit gates and the sharing of parameters in the $R_{X}$, which in turn may affect the performance of QAOA+ to solve UVPs.
+Therefore, QAOA+ is not suitable for such problems.
+For UVPs, we consider performing mixed operators $U(H_{M},\beta)$ on constrained variables and HEA on unconstrained variables.
+To reduce the circuit depth, the phase-separation operators $U(H_{P},\gamma)$ are removed, resulting in the following ansatz
 $|\psi_{p}(\overrightarrow{\gamma},\overrightarrow{\beta})\rangle=U(H_{M},\beta_{p})U_{HEA}(\overrightarrow{\gamma_{p}})\cdots U_{HEA}(\overrightarrow{\gamma_{1}})|x\rangle,$ (3)
-where $\overrightarrow{\gamma_{i}}=(\gamma_{i}^{1},\gamma_{i}^{2},\cdots,\gamma_{i}^{l})$, $i=1,2,\cdots,p$, and $l$ is the number of non-constrained variables.
-$U(H_{M},\beta_{i})$ acts on variables of the constraints, and $U_{HEA}(\overrightarrow{\gamma_{i}})$ acts on non-constrained variables in constrained optimization problems.
-The crucial points of PFS-VQA are the initial state $|x\rangle$, the mixing operators $U(H_{M},\beta)$, the unitary operators $U_{HEA}(\gamma)$, and the phase separator Hamiltonian $H_{P}$.
-The framework of the PFS-VQA is shown in Figure  1.
-Figure 1: The framework of the PFS-VQA.
+where $\overrightarrow{\gamma_{i}}=(\gamma_{i}^{1},\gamma_{i}^{2},\cdots,\gamma_{i}^{l})$, $i=1,2,\cdots,p$, and $l$ is the number of unconstrained variables.
+$U(H_{M},\beta_{i})$ acts on variables of the constraints, and $U_{HEA}(\overrightarrow{\gamma_{i}})$ acts on unconstrained variables in constrained optimization problems.
+The crucial points of VQA-PFS are the initial state $|x\rangle$, the mixing operators $U(H_{M},\beta)$, the unitary operators $U_{HEA}(\gamma)$, and the phase separator Hamiltonian $H_{P}$.
+The framework of the VQA-PFS is shown in Figure  2.
+Figure 2: The framework of the VQA-PFS.
 The variational parameters were optimized on a classical computer, and a quantum computer was used to evaluate the expected value of the objective function.
@@
 where $x_{sol}$ is an optimal solution to the problem.
-III Apply PFS-VQA to solve UFLP
+III Apply VQA-PFS to solve UFLP
 The UFLP is a combinatorial optimization problem that can be transformed into a variant of UVP.
 In this section, we first introduce the definition and the mathematical model of UFLP.
-Then, we solve UFLP using PFS-VQA.
+Then, we solve UFLP using VQA-PFS.
 III.1 UFLP
@@
 The number of binary decision variables is $2mn+n$.
-PFS-VQA combines the advantages of QAOA+ and HEA by performing mixed operators on constrained variables and HEA on unconstrained variables respectively, which searches for a target solution in the feasible space and also decreases the circuit depth.
-However, one of the cores of this algorithm is the construction of mixed Hamiltonian.
+VQA-PFS performs mixed operators on constrained variables and HEA on unconstrained variables, respectively, combining the advantages of QAOA+ and HEA to find the objective solution in the feasible space while reducing the depth of the circuits, and one of the cores is the construction of mixed Hamiltonian.
 To facilitate the creation of the mixed Hamiltonian, we adopt the penalty function approach by making the constraint Eq.
@@
 (13) is $mn+n$.
-Next, we will solve the UFLP via PFS-VQA designed in section  II.
-III.2 PFS-VQA FOR UFLP
-The crucial points of PFS-VQA are the initial state $|x\rangle$, the mixing operators $U(H_{M},\beta)$, the unitary operators $U_{HEA}(\gamma)$, and the phase separator Hamiltonian $H_{P}$ for UFLP.
+Next, we will solve the UFLP via VQA-PFS designed in section  II.
+III.2 VQA-PFS FOR UFLP
+The crucial points of VQA-PFS are the initial state $|x\rangle$, the mixing operators $U(H_{M},\beta)$, the unitary operators $U_{HEA}(\gamma)$, and the phase separator Hamiltonian $H_{P}$ for UFLP.
 For initial state $|x\rangle$, according to Eq.
@@
 where $X$, and $Y$ represents Pauli-$X$ operation, and Pauli-$Y$ operation respectively.
-The unitary operators $U_{HEA}(\gamma)$ of HEA consist of single-qubit rotating gates and entangled gates as shown in Figure  1.
+The unitary operators $U_{HEA}(\gamma)$ of HEA consist of single-qubit rotating gates and entangled gates as shown in Figure  2.
 The phase separator Hamiltonian $H_{P}$ is obtained by replacing binary variables $x$, $y$, $z$ in Eq.
@@
 where $Z$ represents Pauli-$Z$ operation and the subscript represents the qubit of action.
-After the three key points are structured, we apply PFS-VQA to solve the UFLP and perform numerical simulation experiments.
+After the four key points are structured, we apply VQA-PFS to solve the UFLP and perform numerical simulation experiments.
 IV Numerical Simulation
 In this section, we perform numerical experiments using the MindSpore Quantum 51 .
-We study 12 instances for three different problem sizes of the UFLP given in Table  1 to benchmark the performance of PFS-VQA.
+We study 12 instances for three different problem sizes of the UFLP given in Table  1 to benchmark the performance of VQA-PFS.
 Details of the instances are given in Appendix A, corresponding to 10, 14, and 22 qubits, respectively.
@@
 95
-Figure 2: Comparison of the performance of standard QAOA, QAOA+, HEA, and PFS-VQA for the UFLP with 12 instances.
-The circuit depth and the average success probability as a function of the layers are shown for the standard QAOA, QAOA+, HEA, and PFS-VQA.
+Figure 3: Comparison of the performance of standard QAOA, QAOA+, HEA, and VQA-PFS for the UFLP with 12 instances.
+The circuit depth and the average success probability as a function of the layers are shown for the standard QAOA, QAOA+, HEA, and VQA-PFS.
 (a)-(f) The optimization trajectories from optimizer Adam are plotted versus the number of circuit evaluations.
@@
 (a)-(b) $m=2$, $n=2$, (c)-(d) $m=3$, $n=2$, (e)-(f) $m=5$, $n=2$.
-To verify the high efficiency of PFS-VQA, we also apply QAOA, QAOA+, and HEA to UFLP and compared them, giving the Hamiltonians and the corresponding circuits, as specified in Appendix B, C, and D respectively.
+To verify the high efficiency of VQA-PFS, we also apply QAOA, QAOA+, and HEA to UFLP and compared them, giving the Hamiltonians and the corresponding circuits, as specified in Appendix B, C, and D respectively.
 In Fig.
-2, we performed numerical simulations up to $p=6$ using random initialization for 12 instances.
-The circuit depth and average success probability as a function of the layers are shown for the standard QAOA, QAOA+, HEA, and PFS-VQA, respectively.
+3, we performed numerical simulations up to $p=6$ using random initialization for 12 instances.
+The circuit depth and average success probability as a function of the layers are shown for the standard QAOA, QAOA+, HEA, and VQA-PFS, respectively.
 Specifically, for $m=2$, $n=2$, in Fig.
-2 (a) and (b), PFS-VQA has a higher success probability than the other three algorithms, at least 54% higher.
-Meanwhile, PFS-VQA has the same circuit depth as HEA, which is at least 75% lower than QAOA and QAOA+.
+3 (a) and (b), VQA-PFS has a higher success probability than the other three algorithms, at least 54% higher.
+Meanwhile, VQA-PFS has the same circuit depth as HEA, which is at least 75% lower than QAOA and QAOA+.
 For $m=3$, $n=2$, in Fig.
-2 (c) and (d), PFS-VQA has a success probability of at least 58% higher compared to the other three algorithms.
-Moreover, the circuit depth of PFS-VQA is about 26% lower than HEA, and at least 83% lower than QAOA and QAOA+.
+3 (c) and (d), VQA-PFS has a success probability of at least 58% higher compared to the other three algorithms.
+Moreover, the circuit depth of VQA-PFS is about 26% lower than HEA, and at least 83% lower than QAOA and QAOA+.
 For $m=5$, $n=2$, in Fig.
-2 (e) and (f), compared to the other three algorithms, PFS-VQA has at least five times higher success probability.
-And, the circuit depth of PFS-VQA is about 43% lower than HEA, and at least 87% lower than QAOA and QAOA+.
-We find that the success probability of PFS-VQA is higher than the other three algorithms by an increasing percentage as the number of qubits increases.
+3 (e) and (f), compared to the other three algorithms, VQA-PFS has at least five times higher success probability.
+And, the circuit depth of VQA-PFS is about 43% lower than HEA, and at least 87% lower than QAOA and QAOA+.
+We find that the success probability of VQA-PFS is higher than the other three algorithms by an increasing percentage as the number of qubits increases.
 Directly, we conjecture that the performance of this algorithm is less affected by the increase in qubits than QAOA, QAOA+ and HEA.
-The results show that PFS-VQA achieves a significantly higher success probability at low circuit depths compared to the QAOA, QAOA+, and HEA.
+Figure 4: The comparison of the iterations.
+Figure 5: Resource comparison of the standard QAOA, QAOA+, HEA, and VQA-PFS for the UFLP with three different sizes.
+(a), (b), and (c) show the comparison of four algorithms for the 10, 14, and 22 qubits, respectively.
+The results show that VQA-PFS achieves a significantly higher success probability and lower circuit depth compared to the QAOA, QAOA+, and HEA.
 One of the keys to these four algorithms is to construct the phase separator Hamiltonian.
@@
 These two algorithms need to search for the target solutions in the Overall Hilbert Space 31 .
-Differently, PFS-VQA encodes the constraints directly into quantum circuits, limiting the state of the system to the Feasible Space (a subspace of the entire Hilbert space) of the constrained optimization problems, which implies a higher probability of searching a solution compared to these two algorithms.
-Furthermore, for the Eq.
-(13) with non-constrained variables, QAOA+ performs single-qubit rotating gates $R_{X}$ and the phase-separation operators $U(H_{P},\gamma)$ for non-constrained variables (see Appendix C for details).
-Nevertheless, the $U(H_{P},\gamma)$ of QAOA+ introduce lot of CNOT and $R_{Z}$ gates, thereby leading to increased circuit depth.
-Wang et al 34 have already pointed out that the $U(H_{P},\gamma)$ only change the phase, and numerical experiments demonstrated that discarding it hardly decreases the success probability.
-For unconstrained variables, after discarding the phase-separation operator, only single-qubit rotating gates act on the unconstrained variables, lacking entanglement gates, resulting in the inability to traverse enough quantum states, which may lower the success probability of finding a solution.
-Different from QAOA+, PFS-VQA performs the mixed operators and HEA act on different qubits, parallelization can be realized, leading to a lower circuit depth.
-Therefore, PFS-VQA has lower circuit depth and higher success probability at the same low-layer $p$ compared to the standard QAOA, QAOA+, and HEA.
-Figure 3: The comparison of the iterations.
+Differently, VQA-PFS encodes the constraints directly into quantum circuits, limiting the state of the system to the Feasible Space (a subspace of the entire Hilbert space) of the constrained optimization problems, which implies a higher probability of searching for a solution compared to these two algorithms.
+Furthermore, for UVPs, the mixed operators $e^{-i\beta H_{M}}$ of QAOA+ act on the constrained variables, and only the single-qubit rotating gates $R_{X}$ act on the unconstrained variables (see  C for details).
+The representation capability of this circuit is limited by the lack of two-qubit gates and the sharing of parameters in the $R_{X}$, which leads to a lower success probability of QAOA+ for solving the UVPs.
+Different from QAOA+, VQA-PFS removes the phase-separation operators $U(H_{P},\gamma)$ and performs the mixed operators $e^{-i\beta H_{M}}$ and HEA act on different qubits, parallelization can be realized, leading to a lower circuit depth.
+Therefore, VQA-PFS has lower circuit depth and higher success probability at the same low-layer $p$ compared to the QAOA, QAOA+, and HEA.
 To evaluate the convergence speed of our algorithm, we apply QAOA+, QAOA, and HEA to UFLP and compared them.
@@
 4 through numerical simulation.
-The loss value as a function of the iterations with $p=3$ is shown for the standard QAOA, QAOA+, HEA, and PFS-VQA, respectively.
+The loss value as a function of the iterations with $p=3$ is shown for the standard QAOA, QAOA+, HEA, and VQA-PFS, respectively.
 The results show that our algorithm requires only about 125 iterations to converge to the optimal value, which is faster and better convergence compared to the other three algorithms.
-Figure 4: Resource comparison of the standard QAOA, QAOA+, HEA, and PFS-VQA for the UFLP with three different sizes.
-(a), (b), and (c) show the comparison of four algorithms for the 10, 14, and 22 qubits, respectively.
 A crucial question, especially for near-term devices, is how the different algorithms compare with respect to resource overhead.
@@
 However, our algorithm has significantly more parameters than QAOA and QAOA+, and more CNOT gates than HEA.
-In other words, the PFS-VQA obtains better performance in UFLP at the cost of introducing more variational parameters and CNOTs.
+In other words, the VQA-PFS obtains better performance in UFLP at the cost of introducing more variational parameters and CNOTs.
 V CONCLUSION
-In conclusion, we proposed PFS-VQA for solving such UVPs, a combined ansatz that incorporates the advantages of HEA and QAOA+.
-To verify the high efficiency of PFS-VQA, we also applied QAOA, QAOA+, and HEA to UFLP, as well as compare the performance of these four algorithms.
+In conclusion, we proposed VQA-PFS for solving such UVPs, a combined ansatz that incorporates the advantages of HEA and QAOA+.
+To verify the high efficiency of VQA-PFS, we also applied QAOA, QAOA+, and HEA to UFLP, as well as compare the performance of these four algorithms.
 The Hamiltonians and the corresponding circuits are given as specified in Appendix B, C, and D respectively.
-We tested the performance of several instances of UFLP with 10, 14, and 22 qubits, details of the instances are given in Appendix A, finding that PFS-VQA always outperforms the other three algorithms (see Fig.
+We tested the performance of several instances of UFLP with 10, 14, and 22 qubits, details of the instances are given in Appendix A, finding that VQA-PFS always outperforms the other three algorithms (see Fig.
 3 for details).
 Specifically, for $m=2$, $n=2$, in Fig.
-2 (a) and (b), PFS-VQA has a higher success probability than the other three algorithms, at least 54% higher.
-Meanwhile, PFS-VQA has the same circuit depth as HEA, which is at least 75% lower than QAOA and QAOA+.
+3 (a) and (b), VQA-PFS has a higher success probability than the other three algorithms, at least 54% higher.
+Meanwhile, VQA-PFS has the same circuit depth as HEA, which is at least 75% lower than QAOA and QAOA+.
 For $m=3$, $n=2$, in Fig.
-2 (c) and (d), PFS-VQA has a success probability of at least 58% higher compared to the other three algorithms.
-Moreover, the circuit depth of PFS-VQA is about 26% lower than HEA, and at least 83% lower than QAOA and QAOA+.
+3 (c) and (d), VQA-PFS has a success probability of at least 58% higher compared to the other three algorithms.
+Moreover, the circuit depth of VQA-PFS is about 26% lower than HEA, and at least 83% lower than QAOA and QAOA+.
 For $m=5$, $n=2$, in Fig.
-2 (e) and (f), compared to the other three algorithms, PFS-VQA has at least five times higher success probability.
-And, the circuit depth of PFS-VQA is about 43% lower than HEA, and at least 87% lower than QAOA and QAOA+.
+3 (e) and (f), compared to the other three algorithms, VQA-PFS has at least five times higher success probability.
+And, the circuit depth of VQA-PFS is about 43% lower than HEA, and at least 87% lower than QAOA and QAOA+.
 For the largest scale instance, when $p=3$, our algorithm requires only about 125 iterations to converge to the optimal value, which is faster and better convergence compared to the other three algorithms (see Fig.
@@
 5 for details).
-The proposed algorithm is flexible as HEA can be replaced if a more efficient ansatz is available.
-Moreover, our algorithm is general and instructive for solving such UVPs.
+Our algorithm is general and instructive for solving such UVPs.
 Acknowledgements.
-This work is supported by the Beijing Natural Science Foundation (Grant No.
-4222031), the National Natural Science Foundation of China (Grant Nos.
-61976024, 61972048, 62272056), and the 111 Project B21049.
+This work is supported by National Natural Science Foundation of China (Grant Nos. 62371069, 62372048, 62272056), Beijing Natural Science Foundation (Grant No. 4222031) and the 111 Project B21049.
 Appendix A Details of instances
@@
 where $Z$ represents Pauli-$Z$ operation and the subscript represents the qubit of action.
-After the three key points are structured, we apply QAOA to solve the UFLP and perform numerical simulation experiments, and the corresponding 2-layer quantum circuit is shown in Figure  5.
+After the three key points are structured, we apply QAOA to solve the UFLP and perform numerical simulation experiments, and the corresponding 2-layer quantum circuit is shown in Figure  6.
 Figure 6: The overall 2-layer circuit of QAOA.
@@
 where $Z$ represents Pauli-$Z$ operation and the subscript represents the qubit of action.
-After the three key points are structured, we apply QAOA+ to solve the UFLP and perform numerical simulation experiments, and the corresponding 2-layer quantum circuit is shown in Figure  6.
+After the three key points are structured, we apply QAOA+ to solve the UFLP and perform numerical simulation experiments, and the corresponding 2-layer quantum circuit is shown in Figure  7.
 Figure 7: The overall 2-layer circuit of QAOA+.
@@
 (23).
-For Instance 1-Instance 5, the corresponding 2-layer quantum circuit is shown in Figure  7.
+For Instance 1-Instance 5, the corresponding 2-layer quantum circuit is shown in Figure  8.
 Figure 8: The overall 2-layer quantum circuit of HEA.